what is the mass of a single unit cell of gold. gold crystalizes in a face-centered cubic unit cell.

0.54410-८८ का 1.31 110-24 ज ? क 7.१६ 5 4.1 S1024
Answer #1

In face-centered cubic unit cell number of atom = 4

Atomic mass of gold = 197 g/mol

So 1 mol gold has mass = 197 g

1 mol has 6.023 * 1023 atoms

So, 6.023 * 1023 atoms of gold has mass = 197 g

Or, 1 atoms of gold has mass = 197 / 6.023 * 1023 g

Or, 4 atoms of gold has mass = 197 * 4 / 6.023 * 1023 g = 1.31 * 10-21 g

So answer is option a